# von Neumann entropy and quantum version of thermodynamic entropy

### v1

@@ IV.2 Weakly interacting open quantum system
Here we consider an open system which is in contact with the thermal bath. Although the system exchanges heat with the bath, we assume that the interaction with the bath is so weak that the energy spectrum of the system is more or less unaffected by the interaction, and the dynamics of the quantum system can still be described (at-least effectively) by its Hamiltonian.
As before, let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian with |$\ket{n}$ being the eigenket corresponding to eigenvalue $E_{n}$. In the spirit of Gibbs distribution for classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. We can now get the TH entropy, quantified by the quantum Gibbs entropy, as,
$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (11)
Figure 3: The average VN entropy ($\overline{S}_{VN}$) for subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum). Our system size is 16 and subsystem size is 6.
To find the VN entropy of the open system, one takes the following density operator: $\rho=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy. In fact, for overwhelmingly large number of the pure states (of the system-bath combination), it can be quite generally shown that, a weakly interacting open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [29, 30]. As a result, the VN entropy of the open system is,
$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
where the expression of $p_{n}$ is the same as the one appears in the formula for $S_{qGB}$ in Eq. 11. We now see from Eqs. 11 and 12 that the the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium. The detail discussion on the conditions under which this equivalence works can be found in Sec. II.2.
@@ IV.3 Subsystem of a quantum system
As mentioned in the beginning of this section (Sec. IV), it is non-trivial to verify if the VN entropy is equivalent to the TH entropy for a subsystem (of a much larger quantum system in thermal equilibrium). We analyze this possibility here.
We consider a large composite quantum system $AB$ in a state described by the density matrix $\rho$. We also assume that the subsystem $A$ is macroscopically large but much smaller compare to its counterpart $B$. We here show, mainly by the numerical analysis, that the VN entropy of the subsystem $A$ is proportional to its TH entropy.
To first calculate the TH entropy of a subsystem, we again assume that the total system is prepared with an energy lying in $(E,E+\Delta E]$. As discussed earlier, the system will be in a mixed state $\rho$, as appear in Eq. 9. The TH entropy (qBN entropy) of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq. 8, where the $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell. Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $k_{B}\ln D$ or to the $k_{B}\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$. Accordingly,
$S^{sb}_{TH}\propto k_{B}\ln(DOS),$ (13)
@@
Now to calculate the VN entropy of the subsystem, we first need to calculate the reduced density matrix of the subsystem $\rho_{A}=\text{Tr}_{B}(\rho)$, where $\rho$ is the density matrix of the full system, as given in Eq. 9 (with $p_{n}=\frac{1}{D}$). Denoting the reduced density matrix corresponding to the eigenket |$\ket{n}$ by $\rho_{A}^{(n)}$, where $\rho_{A}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$, we have
$\rho_{A}=\frac{1}{D}\sum_{n}\rho_{A}^{(n)}.$ (14)
Since the subsystem is in the mixed state $\rho_{A}$, we now can calculate the VN entropy of the subsystem directly from $\rho_{A}$, i.e,
$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{A}\ln\rho_{A}),$ (15)
where $\rho_{A}$ is given by Eq. 14. For a system which thermalizes (or obeys the eigenstate thermalization hypothesis [21, 22]), we can also have
$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{A}\ln\rho_{A})\approx-\frac{k_{B}}{D}\sum_{n}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)}).$ (16)
This approximate equality works since, as per the hypothesis, the expectation value of a local operator (say, $O$) is almost the same for different eigenkets from a narrow energy window, i.e., $\text{Tr}(\rho_{A}^{(n)}O)\approx\text{Tr}(\rho_{A}^{(m)}O)$. This is possible if the set of eigenvalues of the reduced density matrix $\rho_{A}^{(n)}$ is almost the same for different $n$. This implies that $\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})\approx\text{Tr}(\rho_{A}^{(m)}\ln\rho_{A}^{(m)})$, which guarantees the approximate equality in Eq. 16. If we write $\overline{S}_{VN}=-\frac{k_{B}}{D}\sum_{n}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})$, then $S^{sb}_{VN}\approx\overline{S}_{VN}$ for a thermalized system.
We note that, if the system does not thermalize (or is integrable), then the VN entropy for the individual eigenkets ($-k_{B}\text{Tr}(\rho_{A}^{(n)}\ln\rho_{A}^{(n)})$) can take widely different values. This in fact can be seen in Fig. 2, where the subsystem VN entropy for different eigenkets are plotted against the energy across the spectrum. We here mention that, for all our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
$H=J\left(\sum_{i=1}^{N-1}\vec{S}_{i}\cdot\vec{S}_{i+1}+\Delta_{2}\sum_{i=1}^{N-2}S^{z}_{i}S^{z}_{i+2}\right).$ (17)
In our calculations, we take $J=1$ and the size of the system $N=16$. This model system is integrable when $\Delta_{2}=0$ and is non-integrable (hence thermalizes) when $\Delta_{2}>0$ [31].
To make more sense of the results in Fig. 2, we plot the subsystem’s average VN entropy ($\overline{S}_{VN}$) against $\ln(DOS)$ (separately for the left and right halves of the spectrum). The results are shown in Fig. 3. We find that, for a thermalized system ($\Delta_{2}>0$), its subsystem’s average VN entropy ($\overline{S}_{VN}$) is proportional to $\ln(DOS)$ with a proportionality constant approximately equal to $\frac{N_{sb}}{N}$ (in our calculations, the subsystem size $N_{sb}=6$ and the full system size $N=16$). This proportionality constant takes a different value (a smaller one) if the system is not thermalized. We take $k_{B}=1$ for all our numerical calculations of entropy. Thus our numerical results suggest that,
$S^{sb}_{VN}\approx\overline{S}_{VN}\propto\ln(DOS).$ (18)
From Eqs. 13 and 18, we conclude that, like for isolated and open systems, for a subsystem also the VN entropy is equivalent to the TH entropy (qBN entropy), i.e., $S^{sb}_{VN}\equiv S^{sb}_{TH}$.
@@
It may be worth mentioning here that, the result that the subsystem’s average VN entropy is proportional to system’s $\ln(DOS)$, is earlier reported in [32] in a different context. We also note that, elsewhere [33], the thermodynamic entropy of a subsystem is defined as what we call here $S^{sb}_{VN}$. As discussed in this work, thermodynamic (TH) entropy is appropriately calculated as the logarithm of $DOS$ (since, $S_{TH}\equiv S_{qBN}$ under some general conditions).
@@ V Conclusion
In this work we analyze whether and when the von Neumann (VN) entropy is equal or equivalent to the thermodynamic (TH) entropy. In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly. We show in this study that, for large thermalized isolated or open system, the VN entropy is equivalent to the TH entropy. Next, through numerical analysis, we also show that the VN entropy of a subsystem of a large thermalized system is equivalent to its TH entropy (qBN entropy).

### v2

@@ IV.2 Weakly interacting open quantum system
Here we consider an open system that is in contact with the thermal bath. Although the system exchanges heat with the bath, we assume that the interaction with the bath is so weak that the energy spectrum of the system is more or less unaffected by the interaction, and the dynamics of the quantum system can still be described (at-least effectively) by its Hamiltonian $H$.
As discussed in Sec. II.3, the density matrix that appropriately describes the open system is $\rho_{\beta}=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy. In fact, if the system-bath combination is described by a microcanonical ensemble or for the overwhelmingly large number of pure states (of the system-bath combination), it can be quite generally shown that an open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [11, 10]. Accordingly, the VN entropy of the open system is,
$S_{VN}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (11)
where $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$ with $Q$ being the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here, $\{E_{n}\}$ is the set of eigenvalues of the system Hamiltonian $H$.
Figure 2: The subsystem VN entropy ($S^{sb}_{VN}$) for the individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
As discussed in Sec. II, for the open quantum system, the TH entropy (or equivalently, the qGB entropy) is given by:
$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
where the expression of $p_{n}$ is the same as the one appears in Eq. 11.
Figure 3: The average VN entropy ($\overline{S}^{sb}_{VN}$) of subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum). Here, the system (subsystem) size is 16 (6) and $\Delta_{2}=0.5$.
We now see from Eqs. 11 and 12 that the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium. The detailed discussion on the conditions under which this equivalence works can be found in Sec. II.2.
@@ IV.3 Subsystem of a quantum system
We separately discuss here the question of equivalence between the subsystem entropies when the full system is in a mixed state (described by a microcanonical ensemble) or in a pure state (both typical and atypical).
If the full system is in a mixed state $\rho_{mc}$, described by a microcanonical (mc) ensemble, then the subsystem will be in a Gibbsian distribution $\rho_{\beta}$ [11, 10]. If $\rho_{sb}$ is the reduced density matrix for the subsystem, i.e., $\rho_{sb}=\text{Tr}_{B}(\rho_{mc})$, then $\rho_{sb}\approx\rho_{\beta}=\frac{e^{-\beta H}}{Q}$ with $Q=\text{Tr}_{sb}(e^{-\beta H})$. It may be noted that the first trace (Tr) operation here is over the degrees of freedom of the bath, B (complement of the subsystem in the full system), and the second trace operation is over the degrees of freedom of the subsystem (sb). We also note that, here, $H$ is the Hamiltonian of the subsystem if the interaction between the subsystem and the bath can be ignored, else it is the effective Hamiltonian of the subsystem, determined by $H=-\frac{1}{\beta}\ln(Q\text{Tr}_{B}(\rho_{mc}))$. This equation can be solved self-consistently: initially, we calculate $Q$ from the Hamiltonian of the subsystem (without interaction with the bath) and get the first order approximation of the effective $H$. Next we use this effective $H$ in $Q$ to find the next order approximation of the effective $H$. We continue this till we get a self-consistent result. Fortunately, we do not need to calculate this effective $H$ in the numerical calculations that follow.
If the full system is in a pure state, for the overwhelming majority of pure states (of the total system), the subsystem can be described by a Gabssian distribution $\rho_{\beta}$ [11, 10]. This important result, known as canonical typicality, leaves open the possibility of pure states (of the full system) for which a subsystem is not described by a Gibbsian distribution or for which a subsystem does not thermalize. We note that there are systems whose eigenstates (from a narrow energy window) are such that the reduced density matrices calculated from them are quite different to each other. These systems do not follow the eigenstate thermalization hypothesis or ETH [27, 28, 5].
Above important results immediately imply the following: if the full system is described by a microcanonical ensemble, the VN entropy of a subsystem is equivalent to its TH entropy (qGB entropy). If the full system is in a pure state, for the overwhelming cases (i.e. for the typical states), the VN entropy of a subsystem will be equivalent to its TH entropy (qGB entropy). But for systems for which the ETH does not hold, this equivalence will not work in general (corresponding pure states are atypical). In the following we take a spin system and numerically demonstrate the issue of equivalence between the VN entropy and TH entropy of a subsystem in the different cases mentioned above.
For our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
$H=\sum_{i=1}^{N-1}\vec{S}_{i}\cdot\vec{S}_{i+1}+\Delta_{2}\sum_{i=1}^{N-2}S^{z}_{i}S^{z}_{i+2}.$ (13)
In our calculations, we take the size of the system $N=16$. This model system is integrable when $\Delta_{2}=0$ and is non-integrable (hence thermalizes) when $\Delta_{2}>0$ [35].
To numerically calculate the VN entropy of the subsystem, we first find $\rho_{sb}$ from the state ($\rho$) of the full system: $\rho_{sb}=\text{Tr}_{B}(\rho)$. We then use the formula Eq. 3 to calculate the VN entropy of the subsystem.
To numerically calculate the corresponding TH entropy, consider that the full system is prepared with an energy lying in $(E,E+\Delta E]$. As discussed earlier, the system will be in a mixed state $\rho_{mc}$, as appear in Eq. 8. The TH (or qBN) entropy of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq. 10, where $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell. Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $k_{B}\ln D$ or to the $k_{B}\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$. Accordingly,
$S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto k_{B}\ln(DOS),$ (14)
@@
Now, for different cases, we calculate the VN entropy and the TH entropy of a subsystem to study their equivalence. First, we consider the case when the full system is in an eigenket |$\ket{n}$ of the Hamiltonian of Eq. 13. In this case the subsystem will be represented by the reduced density matrix $\rho_{sb}\equiv\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$. Now using Eq. 3, we can calculate the VN entropy of the subsystem when the full system is in the eigenket |$\ket{n}$:
$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{sb}\ln\rho_{sb}).$ (15)
The values of $S^{sb}_{VN}$ for different eigenkets across the spectrum can be found in Fig. 2 (for $\Delta_{2}=0.0,0.2$ and $0.5$). We also show the corresponding logarithm of the density-of-states (DOS) for the full system. As mentioned earlier, when the system does not thermalize ($\Delta_{2}=0.0$), the individual reduced density matrices are quite different from each other (even within a narrow energy shell) and accordingly, $S^{sb}_{VN}$ takes quite different values as one goes from one eigenket to another within the energy shell. In this case the VN entropy of the subsystem is not, in general, proportional to its qBN entropy, i.e., to $\ln(DOS)$. On the other hand, when the system thermalizes (or the ETH holds), $S^{sb}_{VN}$ takes very close values as one goes from one eigenket to another within an energy shell (see, for example, results for $\Delta_{2}=0.5$). In Fig. 3, we plot the average VN entropy ($\overline{S}^{sb}_{VN}$) of the subsystem against the logarithm of DOS (individually for left and right halves of the spectrum). We take the average of $S^{sb}_{VN}$ over a small energy shell to get better numerical results. The linearity of the plots shows that the VN entropy of a subsystem is equivalent to its TH (or qBN) entropy, i.e., $S^{sb}_{VN}\approx\overline{S}^{sb}_{VN}\equiv S^{sb}_{qBN}$.
Next, we consider the case when the full system is in a mixed state ($\rho_{mc}$), as appears Eq. 8 (with $p_{n}=\frac{1}{D}$). In this case, the subsystem will be in the following mixed state,
$\rho_{sb}=\frac{1}{D}\sum_{n}\rho_{sb}^{(n)},$ (16)
where the reduced density matrix $\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$ is calculated for a single eigenket |$\ket{n}$ of the full system. Here the sum is over the all eigenkets from a small (fixed) energy shell. There are $D$ eigenkets in the shell. Using Eq. 15, we now can calculate the VN entropy of the subsystem directly from $\rho_{sb}$ of Eq. 16. The value of $S^{sb}_{VN}$ across the energy spectrum can be found in Fig. 4.
Figure 4: The subsystem VN entropy ($S^{sb}_{VN}$) when the full system is in the microcanonical density matrix $\rho_{mc}$. Here, the system (subsystem) size is 16 (6).
The plot of $S^{sb}_{VN}$ against $\ln(DOS)$ shows linearity, similar to what we find in Fig. 3. We note that, as expected, we get the said linear plot even when the system is in integrable regime ($\Delta_{2}=0.0$).
@@
It may be worth mentioning here that, the result that the subsystem’s average VN entropy is proportional to system’s $\ln(DOS)$, is earlier reported in [36] in a different context. We also note that, elsewhere [37], the TH entropy of a subsystem is defined as what we call here the VN entropy of a subsystem ($S^{sb}_{VN}$) when the full system is in $\rho_{mc}$. As discussed in this work, the TH entropy of a subsystem is appropriately calculated as the logarithm of $DOS$ (since, $S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS)$ under some general conditions).
@@ V Conclusion
In this work we analyze whether and when the von Neumann (VN) entropy is equal or equivalent to the thermodynamic (TH) entropy. In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly. We find that, when a large isolated system is in a mixed state (described by microcanonical ensemble), its VN entropy is equivalent to its TH entropy. When the system-bath combination is in a mixed state, or in a typical pure state, the VN entropy of the (open) system is equivalent to its TH entropy. This result also holds for a subsystem (of a much larger system) but the analytical demonstration is not straightforward. Through numerical analysis, we show that the VN entropy of a subsystem is equivalent to its TH (qBN) entropy when the full system (subsystem-bath combination) is described by a microcanonical ensemble or when the full system is in a typical pure state. In this work we also numerically demonstrate that, for the states that are atypical (or for which the ETH does not hold), the VN entropy and the TH entropy of a subsystem are not the same.
Acknowledgements.
SS thanks S. Aravinda for useful discussions.